# Contracts with Inspections

### v1

@@ Abstract
In the classical principal-agent hidden-action model, a principal delegates the execution of a costly task to an agent for which he can choose among actions with different costs and different success probabilities to accomplish the task. To incentivize the agent to exert effort, the principal can commit to a contract, which is the amount of payment based on the task’s success. A crucial assumption of this model is that the principal can only base the payment on the outcome but not on the agent’s chosen action.
In this work, we relax the hidden-action assumption and introduce a new model where the principal is allowed to inspect subsets of actions at some cost that depends on the inspected subset. If the principal discovers that the agent did not select the agreed-upon action through the inspection, the principal can withhold payment. This relaxation of the model introduces a broader strategy space for the principal, who now faces a tradeoff between positive incentives (increasing payment) and negative incentives (increasing inspection).
We show how to find the best deterministic incentive-compatible inspection scheme for all monotone inspection cost functions. We then turn to randomized inspection schemes and show that one can efficiently find the best randomized incentive-compatible inspection scheme when the inspection cost function is submodular. We complement this result by showing that it is impossible to efficiently find the optimal randomized inspection scheme for the more general case of XOS inspection cost functions.
@@ 1 Introduction
Contract theory is the field of economics that studies the interaction between two interested parties, a principal and an agent, and seeks to answer the foundational question: “How do we incentivize people to work?” The original model that captures the tension between the two rational individuals involved is called hidden-action principal-agent interaction [28, 29, 25]. The principal aims to incentivize an agent to take some costly action among a set of $n$ potential actions $A$. Each action $a$ that the agent can take incurs a cost $c(a)$ for the agent and results in a distribution over the set of $m$ potential outcomes $O=\{o_{1},\ldots,o_{m}\}$, where each outcome $o_{j}$ is associated with a reward for the principal $r_{j}$. In order to incentivize the agent to exert effort, that will more likely lead to a favorable outcome, the principal can post a payment scheme (or a contract) that maps the set of outcomes to payments for the agent. The goal of the principal is to design a contract that incentivizes the agent to exert effort, which will maximize her own expected utility (reward minus payment). The postulates defining contract design are limited-liability which states that the payments are only one-way from the principal to the agent, and hidden-action, which states that the payments can only depend on the realized outcome and not the action taken by the agent. When outcomes are binary ($m=2$), it is without loss of generality to assume that $r_{1}=0,r_{2}=1$, i.e., success or failure, and each action $a$ is associated with the probability of success $f(a)$. Then, the goal of the principal is to find a single-parameter $\alpha\in[0,1]$ (the payment in case of success), such that the agent is incentivized to take some action $a^{*}$ (i.e., $a^{*}\in\arg\max_{a}\alpha f(a)-c(a)$) with maximum utility for the principal (i.e., maximizing $(1-\alpha)f(a^{*})$).
However, in many real-life scenarios, the principal may be able to examine specific actions performed by the agent. For example, a store owner can send inspectors to check the performance of his salespeople, and a company can install activity monitoring software on the computers of its employees. In such scenarios, the principal could base the payment not only on the outcome but also on the result of the inspection.
@@
In this work, we relax the hidden-action postulate by allowing the principal to inspect the actions taken by the agent, incurring an inspection cost.
@@
However, if the principal is allowed to condition payment on the event of an inspection, she could easily extract the entire welfare by inspecting the welfare-maximizing action $i^{*}=\arg\max_{i}f(i)-c(i)$ with a very low probability $\varepsilon$, and paying the agent $\frac{c(i^{*})}{\varepsilon}$ only if this action is actually inspected. The agent is incentivized to select action $i^{*}$ since the utility from it is non-negative, and all other actions lead to receiving a payment of 00. The principal’s utility is then $f(i^{*})-c(i^{*})-\varepsilon\cdot v(i^{*})$, where $v(i^{*})$ is the cost of inspecting action $i^{*}$. When $\varepsilon\rightarrow 0$, then this utility converges to the maximum social welfare of the problem. This solution concept would be not only trivial but also not very useful in practice since it uses that the agent is willing to accept a payment just in an arbitrarily small probability. Therefore, we focus on natural contracts, where the principal must commit to a payment and inspection distribution in advance and is only permitted to withhold payment if the agent is caught not performing the desired action.
@@
Our setting can be described as follows: We are given a set of actions $A$ the agent can carry out, each associated with a cost $c(a)$, where $c:A\rightarrow\mathbb{R}_{\geq 0}$, a success probability $f(a)$, where $f:A\rightarrow[0,1]$, and an inspection cost function $v:2^{A}\rightarrow\mathbb{R}_{\geq 0}$. The principal then proposes an inspection scheme described by a suggested action $i\in A$, a payment $\alpha$, and an inspection distribution $p$ over subsets of actions. If the agent takes the suggested action $i$, then he is paid $\alpha$ in case of success. If the agent takes some other action $j\neq i$, then, only in the case that neither action $i$ nor action $j$ are inspected, the agent is paid $\alpha$ in case of success. We say that an inspection scheme is incentive-compatible (IC) if the suggested action is the best response for the agent. The principal’s goal is to find the IC inspection scheme that maximizes the utility of the principal.
@@
Regarding the example of a manager hiring an inspector to evaluate the salespeople’s performance, it may not be feasible or cost-effective to conduct inspections consistently. Consequently, the principal may choose to inspect the salespeople with a certain probability, thereby reducing inspection costs. This strategic approach could potentially result in increased utility, as illustrated in the following example.
Example 1.1.
@@
Let us consider the following instance with three actions $A=\{\bot,b,g\}$: $c(\bot)=0,~{}f(\bot)=\frac{1}{10},~{}v(\bot)=1,~{}c(b)=\frac{1}{10},~{}f(b)=\frac{1}{2},~{}v(b)=1,~{}c(g)=\frac{7}{20},~{}f(g)=1,~{}v(g)=\frac{1}{10}$. We assume $v$ to be additive. Under no inspection, we can observe that $\alpha\in[0,\frac{1}{4})$ incentivizes action $\bot$ and offering $\alpha=0$ yields utility $\frac{1}{10}$; $\alpha\in[\frac{1}{4},\frac{1}{2})$ incentivizes action $b$ and offering $\alpha=\frac{1}{4}$ yields utility $\frac{3}{8}>\frac{1}{10}$; $\alpha\in[\frac{1}{2},1]$ incentivizes action $g$ and offering $\alpha=\frac{1}{2}$ yields utility $\frac{1}{2}>\frac{3}{8}$. Thus, the highest utility for the principal is $\frac{1}{2}$ and corresponds to action $g$. If the principal is only allowed to use a deterministic inspection schemes, i.e., always inspect a specific subset of actions, then inspecting any action different from action $g$ yields strictly negative utility to the principal. The principal can thus ask the agent to do action $g$, inspect it, and offer $\alpha=\frac{7}{20}$, leading to a utility of $\frac{13}{20}-\frac{1}{10}=\frac{11}{20}>\frac{1}{2}$, i.e., it is strictly larger than the highest utility under no inspection. Under randomized inspection schemes, the principal can ask the agent to take action $g$, inspect it with probability $\frac{3}{7}$ (and all other actions with probability 00), offer $\alpha=\frac{7}{20}$ (so as to incentivize action $g$ over the other actions), which leads to a utility of $\frac{13}{20}-\frac{3}{70}=\frac{17}{28}>\frac{11}{20}$, i.e., it is larger than in the previous two scenarios.
@@
In our study, we make the assumption that the inspection cost of actions falls within the complement-free hierarchy introduced by Lehmann et al. [33], and widely used in the context of combinatorial auctions [14, 4, 11], combinatorial and multi-agent contracts [16, 18], as well as Prophet Inequalities and pricing [11, 23] to mention a few. This hierarchy encompasses various real-world scenarios, as can be seen in the example with the company using a monitoring software, where it might be cheaper to buy a monitoring software that checks multiple measures simultaneously rather than purchasing a monitoring software for each measure separately. The inspection cost, in this case, exhibits subadditivity or potentially even submodularity.
@@ 1.1 Our Contribution and Techniques
We first show how to design a simple algorithm that finds the optimal deterministic IC inspection scheme in time polynomial in the number of actions for every monotone combinatorial cost function (Theorem 3.1). To find the best deterministic inspection scheme, we make the following observations. Either it is beneficial to inspect the suggested action $i$ and pay the minimal amount to ensure that the action has non-negative utility $\frac{c(i)}{f(i)}$; otherwise, the payment should be such that the agent’s utility from action $i$ is the same as from some other action $j$. This allows us to consider a polynomial number of values of $\alpha$ such that either $\alpha=\frac{c(i)}{f(i)}$ or $\alpha=\frac{c(i)-c(j)}{f(i)-f(j)}$ for some action $j$. For each such $\alpha$, the set of actions that need to be inspected can be calculated easily: They are the actions with strictly larger utility for the agent under contract $\alpha$.
We then shift our focus to randomized inspection schemes. The first of our two main results proves that for the class of submodular inspection cost functions (i.e., decreasing marginals), one can find the optimal randomized inspection scheme in polynomial time. In particular, we show:
Main Result 1 (Theorem 4.1): One can efficiently find the optimal randomized IC inspection scheme for instances with submodular inspection cost functions, given value query access to the inspection cost function.
@@
Main Result 2 (Theorem 5.1): Finding the best randomized IC inspection scheme for instances with XOS inspection cost functions, requires an exponential number of value (or demand) queries.
To demonstrate this hardness result, we carefully construct a family of instances with the same values of $A,f,c$, but with different XOS inspection cost functions parameterized by a set $T$. For each inspection cost function in the family, we show that the optimal randomized inspection scheme is a solution to a specific form of optimization problem that always admits a unique optimal solution. Distinguishing from the inspection cost function what $T$ is, requires an exponential number of value and demand queries. Given the knowledge of $T$, one can easily find the unique optimal inspection scheme, and from the unique optimal inspection scheme, one can easily derive the parameter $T$ of the instance.
Moreover, in Appendix A, we extend Example 1.1 and show that the gap between the optimal deterministic inspection scheme and optimal randomized inspection scheme, can be as large as $\Omega(n)$. Since the utility of the principal is bounded by the welfare achievable in the non-strategic version of the problem, the gap is known to also be at most $O(n)$. Lastly, in Appendix B we discuss the gaps in the principal’s utility between IC inspection schemes, and non-IC inspection schemes for both the deterministic and randomized cases. In particular, we show that there is no gap in the case of deterministic inspection schemes, but there are cases that admit a gap for randomized inspection schemes.
@@ 1.2 Related Work
After the seminal work of Grossman and Hart [25], who initiated the field of “Contract Theory”, and that of Holmstrom and Milgrom [28, 29], recent years have witnessed a surge of interest in algorithmic questions related to this model [15, 16, 17, 18]. Dütting et al. [15, 17] are among the first to model the problem from an algorithmic viewpoint. Their results prove that it is possible to achieve optimality in the worst case through linear contracts, contracts that pay the agent proportionally to the reward associated with an outcome. Numerous other studies have indicated robust optimality of linear contracts in non-Bayesian settings, particularly in the context of max-min criteria [7, 35, 34, 12]. Recently, Kambhampati [30] has shown how, in certain scenarios, randomized contracts can enhance principal’s reward under max-min criteria.
@@
In the context of inspections, Fallah and Jordan [22] study the problem of a single principal interacting with multiple agents: In particular, the principal aims to encourage safety-compliant actions to maximize utility by utilizing payments and a constrained budget for random inspections. Finally, Ball and Knoepfle [6] have investigated the problem of determining the optimal timing for inspections in a time-dependent setting. Focusing on scenarios where agents exhibit impatience over time, the study explores the ideal moments for conducting inspections.
2 Model
We consider a setting of a single principal and a single agent. The agent can choose an action $a\in A$, each is associated with a cost $c(a)\in\mathbb{R}_{\geq 0}$ and a success probability $f(a)\in[0,1]$. We assume that the agent has a null action $\bot\in A$ for which $c(\bot)=0$ which enforces the game to be individually rational for him, and that there are overall $n$ actions (i.e., $|A|=n$). The null action can be viewed as allowing the agent to opt out of the suggested contract. The principal can inspect the actions taken by the agent and pay for the inspection cost. If the principal inspects action set $S\subseteq A$, it costs her $v(S)$, where $v:2^{A}\rightarrow\mathbb{R}_{\geq 0}$ is a combinatorial function which is normalized (i.e., $v(\emptyset)=0$) and monotone (i.e., for every $S\subseteq T\subseteq A$ it holds that $v(S)\leq v(T)$). An instance of our setting is then defined by a quadruplet $(A,c,f,v)$. The principal’s task is to design an inspection scheme consisting of three components: (1) a suggested action $a\in A$, (2) a contract $\alpha\in[0,1]$, and (3) an inspection distribution $p$ which is a distribution over sets that the principal commits to inspect (i.e., $\sum_{S\subseteq A}p(S)=1$, and for all $S\subseteq A$, $p(S)\in[0,1]$). An inspection scheme $(a,\alpha,p)$ is deterministic if there exists $S\subseteq A$, such that $p(S)=1$ and for every $T\neq S$, $f(T)=0$. For an action $a\in A$ and an inspection distribution $p$, we denote by $p(a)\overset{\mathrm{def}}{=\mathrel{\mkern-3.0mu}=}\sum_{S:a\in S}p(S)$.

### v2

# Contract Design Beyond Hidden-Actions

@@ Abstract
In the classical principal-agent hidden-action contract model, a principal delegates the execution of a costly task to an agent. In order to complete the task, the agent chooses an action from a set of actions, where each potential action is associated with a cost and a success probability to accomplish the task. To incentivize the agent to exert effort, the principal can commit to a contract, which is the amount of payment based on the task’s success but not on the hidden-action chosen by the agent.
In this work, we study the contract design framework under binary outcomes where we relax the hidden-action assumption. We introduce new models where the principal is allowed to inspect subsets of actions at some cost that depends on the inspected subset. If the principal discovers that the agent did not select the agreed-upon action through the inspection, the principal can withhold payment. This relaxation of the model introduces a broader strategy space for the principal, who now faces a tradeoff between positive incentives (increasing payment) and negative incentives (increasing inspection).
We devise algorithms for finding the best deterministic and randomized incentive-compatible inspection schemes for various assumptions on the inspection cost function. In particular, we show the tractability of the case of submodular inspection cost functions. We complement our results by showing that it is impossible to efficiently find the optimal randomized inspection scheme for the more general case of XOS inspection cost functions, and that there is no PTAS for the case of subadditive inspection cost functions.
@@ 1 Introduction
Contract theory is the field of economics that studies the interaction between two interested parties, a principal and an agent, and seeks to answer the foundational question: “How do we incentivize people to work?” The original model that captures the tension between the two rational individuals involved is called hidden-action principal-agent interaction [29, 30, 26]. The principal aims to incentivize an agent to take some costly action among a set of $n$ potential actions $A$. Each action $a\in A$ that the agent can take incurs a cost $c(a)$ for the agent and results in a distribution over the set of $m$ potential outcomes $O=\{o_{1},\ldots,o_{m}\}$, where each outcome $o_{j}$ is associated with a reward for the principal $r_{j}$. In order to incentivize the agent to exert effort, that will more likely lead to a favorable outcome, the principal can post a payment scheme (or a contract) that maps the set of outcomes to payments for the agent. The goal of the principal is to design a contract that incentivizes the agent to exert effort, which will maximize her own expected utility (reward minus payment). The postulates defining contract design are limited-liability which states that the payments are only one-way from the principal to the agent, and hidden-action, which states that the payments can only depend on the realized outcome and not the action taken by the agent. When outcomes are binary ($m=2$), it is without loss of generality to assume that $r_{1}=0,r_{2}=1$, i.e., success or failure, and each action $a$ is associated with the probability of success $f(a)$. Then, the goal of the principal is to find a single-parameter $\alpha\in[0,1]$ (the payment in case of success), such that the agent is incentivized to take some action $a^{*}$ (i.e., $a^{*}\in\arg\max_{a}\alpha f(a)-c(a)$) with maximum utility for the principal (i.e., maximizing $(1-\alpha)f(a^{*})$).
However, in many real-life scenarios, the principal may be able to examine specific actions performed by the agent. Consider for example a bank (the principal) that partners with a loan officer (the agent) to evaluate and approve loan applications. The loan officer can choose between conducting thorough evaluations, which are time-consuming and costly, or performing quick but superficial reviews, increasing the risk of approving high-risk loans. Since the bank cannot directly observe the evaluation process, it designs a contract that combines per-loan commissions with random inspections of approved loans. There are many potential types of inspections the bank can do, and each combination of them is associated with (potentially combinatorial) cost. If audits reveal insufficient evaluations, penalties or reduced payments are imposed, incentivizing the loan officer to align his actions with the bank’s low-risk policies while balancing inspection costs. Another example is where a store owner can send inspectors to check the performance of his salespeople, and a company can install activity monitoring software on the computers of its employees. Another motivating example inspired by the concurrent work of Fallah and Jordan [23], involves a principal aiming to incentivize a worker to choose an action of the highest safety level among $L$ available levels, with the principal able to verify whether the safety level of the selected action is at least $\ell$. While Fallah and Jordan [23] focused on scenarios with only two safety levels (safe and unsafe), our framework addresses more complex cases such as having multiple levels. In such scenarios, the principal could base the payment not only on the outcome but also on the result of the inspection.
@@
In this work, we relax the hidden-action postulate by allowing the principal to inspect the actions taken by the agent, incurring an inspection cost. We focus on natural contracts, where the principal must commit to a payment and inspection distribution in advance and is only permitted to withhold payment if the agent is caught not performing the desired action. [^1] If the principal is allowed to condition payment on the event of an inspection, she could easily extract the entire welfare by inspecting the welfare-maximizing action $i^{*}=\arg\max_{i}f(i)-c(i)$ with a very low probability $\varepsilon$, and paying the agent $\frac{c(i^{*})}{\varepsilon}$ only if this action is actually inspected. The agent is incentivized to select action $i^{*}$ since the utility from it is non-negative, and all other actions lead to receiving a payment of 0. The principal’s utility is then $f(i^{*})-c(i^{*})-\varepsilon\cdot v(i^{*})$, where $v(i^{*})$ is the cost of inspecting action $i^{*}$. When $\varepsilon\rightarrow 0$, then this utility converges to the maximum social welfare of the instance. This solution concept would be not only mathematically trivial but also not very useful in practice since it uses that the agent is willing to accept a payment just in an arbitrarily small probability and being paid in cases that are fully controlled by the principal.
@@
Our setting can be described as follows: We are given a set of actions $A$ the agent can carry out, each associated with a cost $c(a)$, a success probability $f(a)$, and an inspection cost function $v:2^{A}\rightarrow\mathbb{R}_{\geq 0}$. The principal then proposes an inspection scheme described by a suggested action $i\in A$, a payment $\alpha$, and an inspection distribution $p$ over subsets of actions. If the agent takes the suggested action $i$, then he is paid $\alpha$ in case of success. If the agent takes some other action $j\neq i$, then, only in the case that neither action $i$ nor action $j$ are inspected, the agent is paid $\alpha$ in case of success. This is what we call fully-observable inspection model. We say that an inspection scheme is incentive-compatible (IC) if the suggested action is the best response for the agent. The principal’s goal is to find the IC inspection scheme that maximizes the utility of the principal.
We also extend the inspection model to a setting where the inspection outcome is only partially-observable. Namely, when the principal inspects some set $S\subseteq A$, she only receives as feedback a single bit indicating whether the action $j$ taken by the agent was in $S$ or not. This is in contrast to the fully-observable model where, by inspecting set $S\subseteq A$, the principal knows exactly which action $j\in S$ (if any) the agent has taken.
@@
Regarding the example of a manager hiring an inspector to evaluate the salespeople’s performance, it may not be feasible or cost-effective to always conduct inspections. Consequently, the principal may choose to inspect the salespeople with a certain probability, thereby reducing inspection costs. This strategic approach could potentially result in increased utility, as illustrated in the following example.
Example $1.$
@@
Let us consider the following instance with three actions $A=\{\bot,b,g\}$: $c(\bot)=0,~f(\bot)=\frac{1}{10},~v(\bot)=1,~c(b)=\frac{1}{10},~f(b)=\frac{1}{2},~v(b)=1,~c(g)=\frac{7}{20},~f(g)=1,~v(g)=\frac{1}{10}$. We assume $v$ to be additive. Under no inspection, we can observe that $\alpha\in[0,\frac{1}{4})$ incentivizes action $\bot$ and offering $\alpha=0$ yields utility $\frac{1}{10}$; $\alpha\in[\frac{1}{4},\frac{1}{2})$ incentivizes action $b$ and offering $\alpha=\frac{1}{4}$ yields utility $\frac{3}{8}>\frac{1}{10}$; $\alpha\in[\frac{1}{2},1]$ incentivizes action $g$ and offering $\alpha=\frac{1}{2}$ yields utility $\frac{1}{2}>\frac{3}{8}$. Thus, the highest utility for the principal is $\frac{1}{2}$ and corresponds to action $g$. If the principal is only allowed to use a deterministic inspection schemes, i.e., always inspect a specific subset of actions, then inspecting any action different from action $g$ yields strictly negative utility to the principal. The principal can thus ask the agent to do action $g$, inspect it, and offer $\alpha=\frac{7}{20}$, leading to a utility of $\frac{13}{20}-\frac{1}{10}=\frac{11}{20}>\frac{1}{2}$, i.e., it is strictly larger than the highest utility under no inspection. Under randomized inspection schemes, the principal can ask the agent to take action $g$, inspect it with probability $\frac{3}{7}$ (and all other actions with probability 0), offer $\alpha=\frac{7}{20}$ (so as to incentivize action $g$ over the other actions), which leads to a utility of $\frac{13}{20}-\frac{3}{70}=\frac{17}{28}>\frac{11}{20}$, i.e., it is larger than in the previous two scenarios.
@@
In our study, we make the assumption that the inspection cost of actions falls within the complement-free hierarchy introduced by Lehmann et al. [35], and widely used in the context of combinatorial auctions [14, 4, 11], combinatorial and multi-agent contracts [16, 18], as well as Prophet Inequalities and pricing [11, 24] to mention a few. This hierarchy encompasses various real-world scenarios, as can be seen in the example with the company using a monitoring software, where it might be cheaper to buy a monitoring software that checks multiple measures simultaneously rather than purchasing a monitoring software for each measure separately. The inspection cost, in this case, exhibits subadditivity or potentially even submodularity (see formal definition in Section 2).
@@ 1.1 Our Contribution and Techniques
We first show how to design a simple algorithm that finds the optimal deterministic IC inspection scheme in time polynomial in the number of actions for every monotone combinatorial cost function (Theorem 3.1) in the fully observable inspection model. To find the best deterministic inspection scheme, we make the following observations. Either it is beneficial to inspect the suggested action $i$ and pay the minimal amount to ensure that the action has non-negative utility $\frac{c(i)}{f(i)}$; otherwise, the payment should be such that the agent’s utility from action $i$ is the same as from some other action $j$. This allows us to consider a polynomial number of values of $\alpha$ such that either $\alpha=\frac{c(i)}{f(i)}$ or $\alpha=\frac{c(i)-c(j)}{f(i)-f(j)}$ for some action $j$. For each such $\alpha$, the set of actions that need to be inspected can be calculated easily: They are the actions with strictly larger utility for the agent under contract $\alpha$.
We then shift our focus to randomized inspection schemes. Our first main result proves that for the class of submodular inspection cost functions (i.e., decreasing marginals), one can find the optimal randomized inspection scheme in polynomial time. In particular, we show:
Main Result 1 (Theorem 4.1): In the fully-observable inspection model, one can efficiently find the optimal randomized IC inspection scheme for instances with submodular inspection cost functions, given value query access to the inspection cost function.
@@
Main Result 2 (Theorem 5.1): In the fully-observable inspection model, finding the best randomized IC inspection scheme for instances with XOS inspection cost functions, requires an exponential number of value (or demand) queries.
To demonstrate this hardness result, we carefully construct a family of instances with the same values of $A,f,c$, but with different XOS inspection cost functions parameterized by a set $T$. For each inspection cost function in the family, we show that the optimal randomized inspection scheme is a solution to a specific form of optimization problem that always admits a unique optimal solution. Distinguishing from the inspection cost function what $T$ is, requires an exponential number of value and demand queries. Given the knowledge of $T$, one can easily find the unique optimal inspection scheme, and from the unique optimal inspection scheme, one can easily derive the parameter $T$ of the instance. We, furthermore, show an stronger inapproximability result for the broader class of subadditive inspection costs:
Main Result 3 (Theorem 6.1): In the fully-observable inspection model, finding a $\nicefrac{{22}}{{21}}$-approximation to the best randomized IC inspection scheme for instances with subadditive inspection cost functions, requires an exponential number of value (or demand) queries.
In Section 7, we extend our results in the partially-observable inspection model. We first show that this model generalizes the (single-agent) model that was considered concurrently and independently in the work [^2] We note that Fallah and Jordan [23] also considered settings with multiple agents which is not captured by our model. of Fallah and Jordan [23]. We also show that, in contrast to the fully-observable inspection model, deterministic inspection schemes can only be found efficiently for submodular functions (see Theorem 7.1) and not all set functions (see Proposition 7.2). In addition, we show that the problem of finding a randomized IC inspection schemes in the partially-observable inspection model admits a convexity property, which allows approximating it:
Main Result 4 (Theorem 7.3): In the partially-observable inspection model, one can efficiently find an additive $\varepsilon$-approximation to the optimal randomized IC inspection scheme for instances with submodular inspection cost functions, given value query access to the inspection cost function.
Moreover, in Appendix C, we extend Example 1.1 and show that the gap between the optimal deterministic inspection scheme and optimal randomized inspection scheme, can be as large as $\Omega(n)$. Since the utility of the principal is bounded by the welfare achievable in the non-strategic version of the problem, the gap is known to also be at most $O(n)$. Lastly, in Appendix B we discuss the gaps in the principal’s utility between IC inspection schemes, and non-IC inspection schemes for both the deterministic and randomized cases. In particular, we show that there is no gap in the case of deterministic inspection schemes, but there are cases that admit a gap for randomized inspection schemes.
@@ 1.2 Related Work
After the seminal work of Grossman and Hart [26], who initiated the field of “Contract Theory”, and that of Holmstrom and Milgrom [29, 30], recent years have witnessed a surge of interest in algorithmic questions related to this model [15, 16, 17, 18]. Dütting et al. [15, 17] are among the first to model the problem from an algorithmic viewpoint. Their results prove that it is possible to achieve optimality in the worst case through linear contracts, contracts that pay the agent proportionally to the reward associated with an outcome. Numerous other studies have indicated robust optimality of linear contracts in non-Bayesian settings, particularly in the context of max-min criteria [7, 37, 36, 12]. Recently, Kambhampati [32] has shown how, in certain scenarios, randomized contracts can enhance principal’s reward under max-min criteria. For more recent developments on algorithmic contract design, see [20].
@@
In the context of inspections, the recent independent and concurrent work of Fallah and Jordan [23] study the problem of a single principal interacting with multiple agents: In particular, the principal aims to encourage safety-compliant actions to maximize utility by utilizing payments and a constrained budget for random inspections. Finally, Ball and Knoepfle [6] have investigated the problem of determining the optimal timing for inspections in a time-dependent setting. Focusing on scenarios where agents exhibit impatience over time, the study explores the ideal moments for conducting inspections.
2 Fully-Observable Inspection Model
We consider a setting of a single principal and a single agent with binary outcomes. The agent can select an action $a\in A$, each is associated with a cost $c(a)\in\mathbb{R}_{\geq 0}$ and a success probability $f(a)\in[0,1]$. If the agent selects action $i$, the task is successful with probability $f(a)$, which gives the principal a reward (which we normalize without loss of generality to $1$). We assume that the agent has a null action $\bot\in A$ for which $c(\bot)=0$ which enforces the game to be individually rational for the agent, and that there are overall $n$ actions (unless stated otherwise). The null action can be viewed as allowing the agent to opt out of the suggested contract. The principal can inspect whether some actions were taken by the agent and pay for the inspection cost. If the principal inspects action set $S\subseteq A$, then she gets to observe whether any action $j\in S$ has been taken by the agent. In Section 7, we analyze different inspection model where the principal only observes whether some action $j\in S$ has been taken by the agent. The inspection costs the principal $v(S)$, where $v:2^{A}\rightarrow\mathbb{R}_{\geq 0}$ is a combinatorial function which is normalized (i.e., $v(\emptyset)=0$) and monotone (i.e., for every $S\subseteq T\subseteq A$ it holds that $v(S)\leq v(T)$). An instance of our setting is then defined by a quadruplet $(A,c,f,v)$. The principal’s task is to design an inspection scheme consisting of three components: (1) a suggested action $a\in A$, (2) a contract $\alpha\in[0,1]$, and (3) an inspection distribution $p$ which is a distribution over sets that the principal commits to inspect (i.e., $\sum_{S\subseteq A}p(S)=1$, and for all $S\subseteq A$, $p(S)\in[0,1]$). An inspection scheme $(a,\alpha,p)$ is deterministic if there exists $S\subseteq A$, such that $p(S)=1$ and for every $T\neq S$, $p(T)=0$. For an action $a\in A$ and an inspection distribution $p$, we denote by $p(a)\overset{\mathrm{def}}{=\mathrel{\mkern-3.0mu}=}\sum_{S:a\in S}p(S)$.